How can we change two body system into one.


1
Expert's answer
2021-02-15T00:54:46-0500

Suppose we have a two body system (body 1, body 2) with a potential depending only on their relative distance. The total energy is :

"E = \frac{m_1 v_1^2}{2} + \frac{m_2 v_2^2}{2} + V(r_1-r_2)"

and thus we introduce the relative position "r=r_1-r_2", position of a center of mass "R=\frac{m_1 r_1 + m_2 r_2}{m_1+m_2}" and we reformulate the positions "r_1, r_2" with these new variables :

"r_1 = R+\frac{m_2}{m_1+m_2} r"

"r_2 = R - \frac{m_1}{m_1+m_2}r"

And now the velocities :

"v_1 = v_{MC} + \frac{m_2}{m_1+m_2} v"

"v_2 = v_{MC} - \frac{m_1}{m_1+m_2} v"

Now by reformulating the energy in these new variables we get :

"E =\frac{(m_1+m_2)v_{MC}^2}{2} + \frac{1}{2} \frac{m_1m_2}{m_1+m_2} v^2+V(r)"

We see that the potential energy is independent from R and thus "v_{MC}" is conserved (which is coherent with the physical explanation - the system is isolated and thus the total momentum is conserved). We can place ourselves in a inertial frame of reference where "v_{MC}=0". Now by introducing "\mu=\frac{m_1 m_2}{m_1 + m_2}" - reduced mass, we get an energy of a one body in a central field :

"E = \frac{\mu v^2}{2} +V(r)"

And now we can solve the problem completely in this formalism.